OpenStudy (anonymous):

how to simplify: 9a-3/8a divided by 3a-1/24

OpenStudy (anonymous):

\[\frac{\frac{9a-3}{8a}}{\frac{3a-1}{24}}\] Let's rewrite that for a second: \[X=\frac{9a-3}{8a}\] So now we just have: \[\frac{X}{\frac{3a-1}{24}}\] Much easier to read :) Ok, so a couple of things: \[1 = \frac{2}{2}=\frac{3}{3}=... = \frac{Y}{Y}=\text{ or even } = \frac{\frac{1}{Y}}{\frac{1}{Y}}\] So, to get rid of the fraction on the bottom, multiply the top and bottom by the inverse of the fraction.